User Tools

Site Tools


lbaops:lbadec2018:v252bpholog

Differences

This shows you the differences between two versions of the page.

Link to this comparison view

Both sides previous revision Previous revision
lbaops:lbadec2018:v252bpholog [2019/02/04 16:58]
breid removed
— (current)
Line 1: Line 1:
-Telescope only joined at ca. 18:45 UT. Around 23:00 UT receiver temperature increased from 70 K to 130 K due to the possible leakage of the cooling fluid. After pump was turned on for cooling, the temperature dropped back to 100 K by 00:30 UT and stabilized.